# Bounds on $k$-hash distances and rates of linear codes

### v2

@@ V A Related Open Problem in Zero-Error Capacity
As mentioned in the introduction, the problem of perfect hashing can be alternatively described as a problem of zero-error capacity under list decoding for a certain class of discrete memoryless channels. In particular, $(q,k)$-hash codes are zero-error codes with decoding list-size $k-1$ for symmetric channels with $q$ inputs and $\binom{q}{k-1}$ outputs, each one reachable from $k-1$ inputs. For example, for the channels in Figure 2, a code is error-free under list-decoding with list-size $2$ if any three distinct codewords have, in at least one coordinate, three distinct symbols. Thus, zero-error codes are, respectively, $(3,3)$-hash (trifferent) and $(4,3)$-hash codes. In this section we want to bring the attention to the possible use of the method presented in Section III in the more general context of bounding zero-error capacities. In particular, we consider a specific case that in our opinion represents a significant example for more general studies in these kind of problems. While the problem seems to be similar to the one studied in the previous sections, and the method can in principle also be applied in a similar way, it does not give useful results. So, this section is in a certain sense a story of failure, but we believe the problem raised is of sufficient interest to be still discussed in relation to the present work.
@@
Figure 2: Channel with a positive zero-error list-decoding capacity with list-size $2$; no output is compatible with more than two inputs. Zero-error codes for these two channels are, respectively, $(3,3)$-hash (trifferent) and $(4,3)$-hash codes.
@@
Figure 3: On the left, a $5$-input typewriter channel. On the right, its confusability graph. Zero-error codes under list-decoding with list-size $2$ correspond in this case to triangle-free subgraphs of the strong power of the pentagon.
The specific channel we consider is the famous example, studied by Shannon [27] and Lovász [28], of the typewriter channel with $5$ inputs, shown in Figure 3. Zero-error codes under ordinary (list-size 1) decoding correspond to independent sets in the strong power of the confusability graph of the channel, which in this case is a pentagon. For general channels, with larger list-sizes the confusability graph is no more relevant and one has to consider hypergraphs. Typewriter channels, though, constitute a special case for which the confusability graph still contains all important information, leading to an interesting variation of the so called “graph capacity” problem [28].
@@
Consider the channel of Figure 3 and its confusability graph. Under ordinary decoding (list-size 1), zero-error codes of length $n$ for this channel can be interpreted as independent sets in the strong power $C_{5}^{n}$ of the pentagon [28]. Shannon observed that for $n=2$ there exists a zero-error code with $5$ codewords $\{00,12,24,31,42\}$. Note that this code is linear as a code over ${\mathbb{F}}_{5}$. This trivially implies that, for even $n$, there exists a linear zero-error code with $5^{n/2}$ codewords. Lovász [28] proved that these codes are of maximal size, thus determining that the so-called “graph capacity” of the pentagon is $1/2$. So, under ordinary decoding, the zero-error capacity is achieved by linear codes. Consider then zero-error codes under list-decoding with list-size $2$ for this channel. The requirement that no output sequence be compatible with three distinct codewords can be here rephrased (for this special case) as the property that among any three codewords, at least two of them are not confusable. Thus, with list-size 2, zero-error codes correspond to triangle-free subgraphs of the powers of the pentagon. So, on the one hand this problem suggests an interesting extension of the notion of “graph capacity”. On the other hand, the channel can be considered as a natural way of extending the first channel of Figure 2, so in a sense this represents another possible way of extending the trifference problem [^4] The first significant one using so-called “typewriter” channels, since for an even number of inputs the zero-error capacity remains the same for any list-size. as opposed to the second channel of Figure 2.
@@
First observe that standard zero-error codes for list-size 1 also clearly work for list-size 2, so we known that the capacity is at least $1/2$. To the best of our knowledge, no better lower bound is known. For the upper bounds, Lovász method cannot be easily adapted to this situation, and the best known (to our knowledge) upper bound is $\log_{5}(5/2)$. This comes from the same simple packing argument used to derive (1), and it also holds for any larger list-size. So, the best known lower bound is achieved by linear codes, and the best known upper bound holds not just for list-size $2$ but for any list-size. In light of the results of [18][19], it is rather natural to ask whether one can at least improve the upper bound for linear codes. In principle, one can combine the results in [29] with the methods in this paper. This points out how bounds on the minimum distance of codes at rates above graph capacity as studied in [29] can be useful for other problems not considered there. In practice, unfortunately, the method does not lead to an improvement on the $\log_{5}(5/2)$ bound. Still, the problem which emerges is rather interesting and it deserves in our opinion being discussed here and further studied.
@@
Define, as in [29], the distance between two codewords in ${\mathbb{F}}_{5}$ for the channel of Figure 3 as their Hamming distance if they are confusable, and infinite otherwise. Thus, zero-error codes have infinite minimum distance, while the minimum distance of any code at rate $R>1/2$ is finite. Given a linear code of length $n$ and rate $R=m/n>1/2$, let $x_{1}$ be a codeword at minimum distance $d=\delta n$ from 00. Assume again $x_{1}$ contains $1$’s in the first $d$ positions and zeros in the remaining $n-d$ ones. If the code is list-decodable with list-size $2$, any other codeword $x_{2}$ must either be distinguishable from 00 or from $x_{1}$. So, either it contains a symbol $2,3$ or $4$ in the first $d$ coordinates, or a symbol $2$ or $3$ in the last $n-d$. Using Lemma 1 as done in Section II-B, we deduce that
$3d+2(n-d)+1\geq 4m\,,$
@@
$\delta\leq 4R-2\,.$ (17)
This can be combined with any rate-distance converse bound as was done in Section II-B using the Plotkin or the Linear Programming bound. However, what we need in this case is a bound for this unusual distance which takes infinite values for non-confusable symbols. This type of bounds were studied in [29]. According to Theorem 2 there, the rate and minimum distance of any code for the considered channel satisfy
$R\leq\frac{1}{2}+\frac{1}{2}R_{LP1}(\sqrt{5},\delta)\,.$
where $R_{LP1}(q,\delta)$ is defined in (11). Combining this with (17) gives an upper bound on the rate which is determined by solving
$R=\frac{1}{2}+\frac{1}{2}R_{LP1}(\sqrt{5},4R-2).$
@@
Unfortunately, this gives $R\leq 0.593$, and so no improvement is obtained over the simple upper bound $\log_{5}(5/2)\approx 0.5693$ (see Figure 4). What is particularly interesting for this problem is that by looking at the bounds on $\delta$ derived in [29], one finds that in order to obtain even a tiny improvement over $\log_{5}(5/2)$, one would need to improve the upper bounds on $\delta$ in [29] to essentially almost match the best lower bounds presented there, which also holds for linear codes. The margin of improvement would be of only about 0.001, see Figure 4. So, in this case even for linear codes there is little hope to improve the trivial $\log_{5}(5/2)$ bound by using the methods which have been proved effective for the trifference problem, and more generally for $(q,k)$-hash codes. Hence, unless the true achievable rate is really $\log_{5}(5/2)$ (or very close to) even for linear codes, or this problem shows a clear limitation of the considered method. It thus seems to be one worth considering as a challenging next step to attempt.

### v3

@@ IV A Related Open Problem in Zero-Error Capacity
In this section we want to bring the attention to the possible use of the method presented in Section III in the more general context of bounding zero-error capacities. In particular, we consider a specific case that in our opinion represents a significant example for more general studies in these kind of problems. While the problem seems to be similar to the one studied in the previous sections, and the method can in principle also be applied in a similar way, it does not give useful results. So, this section is in a certain sense a story of failure, but we believe the problem raised is of sufficient interest to be still discussed in relation to the present work.
@@
Figure 4: Channel with a positive zero-error list-decoding capacity with list-size 22; no output is compatible with more than two inputs. Zero-error codes for these two channels are, respectively, $(3,3)$-hash (trifferent) and $(4,3)$-hash codes.
@@
Figure 5: On the left, a 55-input typewriter channel. On the right, its confusability graph. Zero-error codes under list-decoding with list-size 22 correspond in this case to triangle-free subgraphs of the strong power of the pentagon.
@@
The specific channel we consider is the famous example, studied by Shannon [29] and Lovász [30], of the typewriter channel with $5$ inputs, shown in Figure 5. Zero-error codes under ordinary (list-size 1) decoding correspond to independent sets in the strong power of the confusability graph of the channel, which in this case is a pentagon. For general channels, with larger list-sizes the confusability graph is no longer relevant and one has to consider hypergraphs. Typewriter channels, though, constitute a special case for which the confusability graph still contains all important information, leading to an interesting variation of the so called “graph capacity” problem [30]. Consider the channel of Figure 5 and its confusability graph. Under ordinary decoding (list-size 1), zero-error codes of length $n$ for this channel can be interpreted as independent sets in the strong power $C_{5}^{n}$ of the pentagon [30]. Shannon observed that for $n=2$ there exists a zero-error code with $5$ codewords $\{00,12,24,31,42\}$. Note that this code is linear as a code over ${\mathbb{F}}_{5}$. This trivially implies that, for even $n$, there exists a linear zero-error code with $5^{n/2}$ codewords. Lovász [30] proved that these codes are of maximal size, thus determining that the so-called “graph capacity” of the pentagon is $1/2$. So, under ordinary decoding, the zero-error capacity is achieved by linear codes. Consider then zero-error codes under list-decoding with list-size $L=2$ for this channel. The requirement that no output sequence be compatible with three distinct codewords can be here rephrased (for this special case) as the property that among any three codewords, at least two of them are not confusable. Thus, with list-size $L=2$, zero-error codes correspond to triangle-free subgraphs of the powers of the pentagon. So, on the one hand this problem suggests an interesting extension of the notion of “graph capacity”. On the other hand, the channel can be considered as a natural way of extending the first channel of Figure 4, so in a sense this represents another possible way of extending the trifference problem [^4] The first significant one using so-called “typewriter” channels, since for an even number of inputs the zero-error capacity remains the same for any list-size. as opposed to the second channel of Figure 4.
@@
First observe that standard zero-error codes for list-size 1 also clearly work for list-size 2, so we known that the capacity is at least $1/2$. To the best of our knowledge, no better lower bound is known. For the upper bounds, Lovász method cannot be easily adapted to this situation, and the best known (to our knowledge) upper bound is $\log_{5}(5/2)$. This comes from the same simple packing argument used to derive (1), and it also holds for any larger list-size. So, the best known lower bound is achieved by linear codes, and the best known upper bound holds not just for list-size $2$ but for any list-size. In light of the results discussed above for the trifference problem, it is rather natural to ask whether one can at least improve the upper bound for linear codes. In principle, one can combine the results in [31] with the methods in this paper. This points out how bounds on the minimum distance of codes at rates above graph capacity as studied in [31] can be useful for other problems not considered there. In practice, unfortunately, the method does not lead to an improvement on the $\log_{5}(5/2)$ bound. Still, the problem which emerges is rather interesting and it deserves in our opinion being discussed here and further studied.
@@
Define, as in [31], the distance between two codewords in ${\mathbb{F}}_{5}$ for the channel of Figure 5 as their Hamming distance if they are confusable, and infinite otherwise. Thus, zero-error codes have infinite minimum distance, while the minimum distance of any code at rate $R>1/2$ is finite. Given a linear code of length $n$ and rate $R=m/n>1/2$, let $x_{1}$ be a codeword at minimum distance $d_{2}=\delta_{2}n$ from 0. Assume again $x_{1}$ contains $1$’s in the first $d$ positions and zeros in the remaining $n-d_{2}$ ones. If the code is list-decodable with list-size $2$, any other codeword $x_{2}$ must either be distinguishable from 0 or from $x_{1}$. So, either it contains a symbol $2,3$ or $4$ in the first $d_{2}$ coordinates, or a symbol $2$ or $3$ in the last $n-d_{2}$. Using Lemma 1 as done in Section II-B, we deduce that
$3d_{2}+2(n-d_{2})+1\geq 4m\,$,
@@
$R\leq\frac{1}{4}\delta_{2}+\frac{1}{2}+o(1)\,.$ (24)
This can be combined with any rate-distance converse bound as was done in Section II-B using the Plotkin or the Linear Programming bound. However, what we need in this case is a bound for this unusual distance which takes infinite values for non-confusable symbols. This type of bounds was studied in [31]. According to Theorem 2 therein, the rate and minimum distance of any code for the considered channel satisfy
$R\leq\frac{1}{2}+\frac{1}{2}R_{LP1}(\sqrt{5},\delta_{2})\,.$
where $R_{LP1}(q,\delta)$ is defined in (19). Combining this with (24) gives an upper bound on the rate of the form $R\leq\frac{1}{4}\delta^{*}+\frac{1}{2}$ where $\delta^{*}$ solves the equation $\delta=2R_{LP1}(\sqrt{5},\delta)$.
@@
Unfortunately, this gives $R\leq 0.593$, and so no improvement is obtained over the simple upper bound $\log_{5}(5/2)\approx 0.5693$ (see Figure 6). What is particularly interesting for this problem is that by looking at the bounds on $\delta_{2}$ derived in [31], one finds that in order to obtain even a tiny improvement over $\log_{5}(5/2)$, one would need to improve the upper bounds on $\delta_{2}$ in [31] to essentially almost match the best lower bounds presented there, which also hold for linear codes. The margin of improvement would be of only about 0.001, see Figure 6. So, in this case even for linear codes there is little hope to improve the trivial $\log_{5}(5/2)$ bound by using the methods which have been proved effective for the trifference problem, and more generally for $(q,k)$-hash codes. Hence, either the true achievable rate is really $\log_{5}(5/2)$ (or very close to) even for linear codes, or this problem shows a clear limitation of the considered method. It thus seems to be one worth considering as a challenging next step.